OpenStudy (anonymous):

Use de Moivre’s Formula to find integers A, B, C such that cos 4θ = Acos^4 θ +Bcos^2 θ +C.

ganeshie8 (ganeshie8):

\[\cos (4\theta) = \mathbb{Real} ~ e^{i 4\theta} = \mathbb{Real} ~ (\cos \theta + i \sin \theta)^4 \]

ganeshie8 (ganeshie8):

expand it using binomial thm, take the real part and compare coefficients

OpenStudy (anonymous):

What do you mean by take the real part?

ganeshie8 (ganeshie8):

take the terms that dont have "\(i\)" in them

ganeshie8 (ganeshie8):

for example : Real part of \(a + ib\) is \(a\)

ganeshie8 (ganeshie8):

\[\cos (4\theta) = \mathbb{Real} ~ e^{i 4\theta} = \mathbb{Real} ~ (\cos \theta + i \sin \theta)^4 \\ =\mathbb{Real} ~ \cos^4 \theta + 4 \cos^3 \theta (i \sin \theta ) + 6 \cos^2 \theta (i \sin \theta)^2 + 4 \cos \theta (i \sin \theta )^3 + (i \sin \theta)^4 \\ = \mathbb{Real} ~ \cos^4 \theta + 4i \cos^3 \theta\sin \theta - 6 \cos^2 \theta sin^2 \theta - 4i \cos \theta \sin^3 \theta + \sin^4 \theta \]

ganeshie8 (ganeshie8):

take the real part of above ^

OpenStudy (anonymous):

Thank you :)

ganeshie8 (ganeshie8):

np :) u will need to change everything into cosines and compare

ganeshie8 (ganeshie8):

\[\cos (4\theta) = \mathbb{Real} ~ e^{i 4\theta} = \mathbb{Real} ~ (\cos \theta + i \sin \theta)^4 \\ =\mathbb{Real} ~ \cos^4 \theta + 4 \cos^3 \theta (i \sin \theta ) + 6 \cos^2 \theta (i \sin \theta)^2 + 4 \cos \theta (i \sin \theta )^3 + (i \sin \theta)^4 \\ = \mathbb{Real} ~ \cos^4 \theta + 4i \cos^3 \theta\sin \theta - 6 \cos^2 \theta \sin^2 \theta - 4i \cos \theta \sin^3 \theta + \sin^4 \theta \\ = \cos^4 \theta - 6 \cos^2 \theta \sin^2 \theta + \sin^4 \theta \]

ganeshie8 (ganeshie8):

so \[ \cos 4\theta = \cos^4 \theta - 6 \cos^2 \theta \sin^2 \theta + \sin^4 \theta = A\cos^4 \theta +B\cos^2 \theta +C \]

ganeshie8 (ganeshie8):

rewrite sin^2 as 1-cos^2 and get it into the right hand side form, then compare